OpenStudy (anonymous):

Simplify 5x-10/5 x − 10 5x − 2 x − 2 x − 5

OpenStudy (anonymous):

5x-10=5(x-2) so 5x-10/5=x-2

Parth (parthkohli):

5(x - 2) =5x - 10 x - 2 is the answer

OpenStudy (anonymous):

you basically do the oppisite from the last equation you divide from both sides

OpenStudy (anonymous):

@ParthKohli u done it again!!!i was going to answer tht!!!!

OpenStudy (anonymous):

and 1x = x

OpenStudy (anonymous):

i know john he is like so fast!!!!

Parth (parthkohli):

*feels like a sir*

OpenStudy (anonymous):

iits wrong answer buddy 5x-10/5=0 25x-10=0 25x=10 x=10/25 x=2/5

Parth (parthkohli):

@ashwinjohn3 Not in the choices, check it lol.

OpenStudy (anonymous):

So,the answer should be 5x -2

OpenStudy (anonymous):

@ParthKohli now check it man!!!!

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 7 * 8 = 56 }\) \(\Large \color{purple}{\rightarrow 56 \div 8 = 7 }\) Basic example

OpenStudy (anonymous):

parth is crrect dude srry

Parth (parthkohli):

@ashwinjohn3 You don't cancel when you add/subtract

OpenStudy (anonymous):

hey u didnt see the division part...buddy

Parth (parthkohli):

See, When you add/subtract the numerator and divide that, then you can't cancel.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {{56 + 56} \over 56} }\) Can you say that the answer is 56?

OpenStudy (anonymous):

let me write it down neatly 5x-10/5(SEE THERE IS A DIVISION PART)=0 5x-2=0 so 5x-2 is the answer

Parth (parthkohli):

Okay, now substitute x with 2.

OpenStudy (anonymous):

x=2/5

Parth (parthkohli):

5x - 10/5 (10 - 10)/5 0/5 0 x - 2 is the answer and x = 2 2 - 2 = 0 CONFIRMED!

OpenStudy (anonymous):

ok.....thnks